Solve for the Electric field of a point charge q, with a value of 2.9 nC and a distance of 2.4 m


1
Expert's answer
2022-02-28T10:38:08-0500

The electric field of a point charge

"E=k\frac{q}{r^2}=9*10^9*\frac{2.9*10^{-9}}{2.4^2}=4.5\:\rm N/C"